# LVAgent: Long Video Understanding by Multi-Round Dynamical Collaboration of MLLM Agents

### v2

@@ Abstract
Existing Multimodal Large Language Models (MLLMs) encounter significant challenges in modeling the temporal context within long videos. Currently, mainstream Agent-based methods use external tools (e.g., search engine, memory banks, OCR, retrieval models) to assist a single MLLM in answering long video questions. Despite such tool-based support, a solitary MLLM still offers only a partial understanding of long videos, resulting in limited performance. In order to better address long video tasks, we introduce LVAgent, the first framework enabling multi-round dynamic collaboration of MLLM agents in long video understanding. Our methodology consists of four key steps: 1) Selection: We pre-select appropriate agents from the model library to form optimal agent teams based on different tasks. 2) Perception: We design an effective retrieval scheme for long videos, improving the coverage of critical temporal segments while maintaining computational efficiency. 3) Action: Agents answer long video-related questions and exchange reasons. 4) Reflection: We evaluate each agent’s performance in each round of discussion and optimize the agent team for dynamic collaboration. The agents iteratively refine their answers by multi-round dynamical collaboration of MLLM agents. LVAgent is the first agent system method that outperforms all closed-source models (including GPT-4o) and open-source models (including InternVL-2.5 and Qwen2-VL) in the long video understanding tasks. Our LVAgent achieves an accuracy of 80% on four mainstream long video understanding tasks. Notably, on the LongVideoBench dataset, LVAgent improves accuracy by up to 13.3% compared with SOTA.
@@ 1 Introduction
Agent-based approaches [9, 49, 16, 51, 57] hold great potential in long video understanding, because agents can simplify difficult problems and autonomously invoke a variety of tools to extract key information from long videos to assist in long video understanding. However, there is still significant room for improvement in current agent-based methods for long video tasks.
Some methods use CLIP [39] to retrieve key frames for long video understanding [49, 57, 64, 55]. However, CLIP struggles to retrieve long-term temporal information and its pretrained data has a domain gap with long videos [46, 59]. Some approaches use external tools (e.g., memory bank [9, 13], RAG [21, 3, 16]) to aid MLLMs in long video understanding. Even with the support of tools, relying on a single MLLM to answer questions about long-form videos offers only a partial comprehension of the video content, leading to limited performance. In conclusion, there are two main challenges in long video understanding. The first challenge is how to better retrieve the video clip with key information according to the query. The second challenge is how to use multi-agent collaboration to better understand long videos.
@@
  •  ̵Second, LVAgent is a novel multi-round and multi-step dynamical collaboration pipeline of MLLM Agents. LVAgent enables multi MLLM agents to process and integrate long-term temporal context dynamically for more comprehensive and robust long video understanding.
  • Finally, empirical results demonstrate the superior accuracy of our LVAgent. LVAgent achieves an accuracy exceeding 80% across four mainstream long video understanding tasks. It is the first multi-round dynamic multi-agent collaboration pipeline that outperforms all closed-source models (including GPT) and open-source models (including InternVL-2.5 and Qwen2-VL) in the long video understanding tasks. Notably, on the LongVideoBench dataset, LVAgent improves accuracy by up to 14.3% compared with SOTA.
@@ 3 Method
Table 1: Comparison on Mainstream Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs”. We are the first agent-based method to achieve over 80% accuracy on each long video understanding dataset
3.1 Selection
3.2 Perception
$D_{i}=A_{i}(\hat{v},Q,T,O)$ (2)
$K_{i}=A_{i}(\hat{v},Q,T,O)$ (3)
$K_{i}=A_{i}(H_{info},Q,T,O)$ (4)
$Sim(chk_{i})=ASP(f_{i},K_{i})+ASP(f_{i},Q)$ (5)
LongVR dataset for Long Video Retrieval.
To enhance the retrieval ability of ASP-CLIP to better assist with long video tasks, we compile a diverse dataset for fine-tuning the ASP-CLIP model, consisting of 82k video clips from five datasets, as shown in Table 8. This collection covers a wide range of scenes, themes, and video quality, providing a robust foundation for training a video retrieval model. We extract long videos from the datasets of ActivityNet [19], OpenVid-1M [36], ViTT [15], MovieChat-Caption [43], and Youcook2 [67] for the retrieval task, as depicted in Figure 6. First, we crop the videos clips with captions. If a video has been pre-annotated into multiple clips, we take these clips along with their captions as components of our dataset. Second, to better suit the long video scenario, we filter out video clips shorter than 5 seconds or longer than 12 minutes. Only clips of appropriate length and their captions are retained. Finally, considering that the context length of CLIP is limited, to prevent information truncation in the clip-text encoder due to overly long captions, we use Qwen2-VL [47] for caption optimization. Additionally, if the number of tokens in a caption is less than 20, we also filter out this data, as such short captions are insufficient to summarize the content of a relatively long video. After manually screening the captions, the resulting long video dataset has an average video duration of 145.6 seconds. The captions in the dataset have an average length of 71 tokens. In total, we assemble a long video retrieval dataset with 82K samples. The details of dataset constructions and data recipes can be found in the supplementary document. We finetune ASP-CLIP [4] on LongVR and achieved significant performance improvements.
@@ 4.2 Implication Details
4.3 Comparison with SOTA
As shown in Table 1, we compare with the best baselines to our work, including closed-source models [38, 44], open-source MLLMs [3, 24, 5, 47, 61, 40, 63, 22, 28, 20, 7], and agent-based systems [49, 51, 57, 33, 32]. Experimental results show that our method is the first agent-based method to achieve over 80% accuracy across these four benchmarks, outperforming all the open-source and closed-source MLLMs. Especially on LongVideoBench, LVAgent outperforms GPT-4o (SOTA) by 13.3%. On MLVU, LVAgent surpasses GPT-4o and InternVL-2.5 by 19.3% and 8.2%.
@@ 4.3 Comparison with SOTA
We evaluate LVAgent’s efficiency on the VideoMME benchmark, comparing it with state-of-the-art models in inference time and performance (Table 2). LVAgent processes 71.2 frames and 15.1s of inference time. Compared with the current MLLM models that perform best in long video tasks, our LVAgent can achieve results that significantly surpass these MLLMs while using fewer frames and less inference time.
@@ 4.4 Ablation Study
As shown in Table 4, we analyze the impact of agent combinations on LVBench and MLVU tasks using LLaVA-Video-72B (LV-72B), InternVL-2.5-8B (Intern-8B), and InternVL-2.5-78B (Intern-78B).
Individually, LV-72B, Intern-8B, and Intern-78B achieve 68.3%, 67.8%, and 69.0% on LVBench, respectively. However, combining all three agents yields superior performance (80.0% on LVBench and 83.9% on MLVU). This improvement stems from the diversity of agents, enabling the model to leverage their complementary strengths for more robust reasoning in long video understanding tasks.
@@
We investigate the effect of varying the number of maximum reflection rounds one, two, and three on multiple benchmarks, as illustrated in Figure 7. Increasing the number of rounds consistently enhances performance; for instance, accuracy on EgoSchema improves from 80.2% with one round to 82.9% with three rounds, while LongVideoBench sees an increase from 69.0% to 80.0%. This improvement is attributed to the iterative refinement and consensus among agents, where multiple rounds progressively eliminate noisy knowledge and converge on more accurate answers, significantly boosting performance in complex reasoning.
@@
We further explore the impact of different retrieval models used in the Perception process, as shown in Table 5. We compare three retrieval models: CLIP, LongCLIP, and ASP-CLIP pretrained on our collected LongVR dataset. The results indicate that ASP-CLIP consistently outperforms the other models across all benchmarks, owing to its superior temporal alignment capabilities.
@@ 6 Acknowledgement
We present a detailed breakdown of the data distribution and key statistics for LongVR dataset in Table 8. Specifically, ActivityNet-Caption contains 37,421 instances with an average caption length of 15.80 words and an average video duration of 171.44 seconds. OpenVid-1M comprises 30,000 instances, featuring longer captions with an average length of 143.16 words and shorter videos averaging 9.07 seconds. ViTT has 5,086 instances, with captions averaging 24.98 words and videos lasting 305.08 seconds on average. MovieChat-Caption includes 808 instances, having the longest captions at an average of 143.47 words and videos with an average duration of 457.65 seconds. Youcook2 consists of 8,700 instances, with captions averaging 80.90 words and videos lasting 383.19 seconds. Collectively, the LongVR dataset encompasses a total of 82,015 instances, with an overall average caption length of 71.12 words and an average video duration of 145.62 seconds. This diverse and extensive dataset enables robust training of the ASP-CLIP model, capturing a wide range of video and caption characteristics.
@@
Appendix B Prompts of LVAgent.

### v3

@@ Abstract
Existing MLLMs encounter significant challenges in modeling the temporal context within long videos. Currently, mainstream Agent-based methods use external tools to assist a single MLLM in answering long video questions. Despite such tool-based support, a solitary MLLM still offers only a partial understanding of long videos, resulting in limited performance. In order to better address long video tasks, we introduce LVAgent, the first framework enabling multi-round dynamic collaboration of MLLM agents in long video understanding. Our method consists of four key steps: 1) Selection: We pre-select appropriate agents from the model library to form optimal agent teams based on different tasks. 2) Perception: We design an effective retrieval scheme for long videos, improving the coverage of critical temporal segments while maintaining computational efficiency. 3) Action: Agents answer long video questions and exchange reasons. 4) Reflection: We evaluate each agent’s performance in each round of discussion and optimize the agent team for dynamic collaboration. The agents iteratively refine their answers by multi-round dynamical collaboration of MLLM agents. LVAgent is the first agent system method that outperforms all closed-source models (like GPT-4o) and open-source models (like InternVL-2.5 and Qwen2-VL) in the long video understanding tasks. Our LVAgent achieves an accuracy of 80% on four mainstream long video understanding tasks. Notably, LVAgent improves accuracy by 13.3% on LongVideoBench. Code is available at https://github.com/64327069/LVAgent.
@@ 1 Introduction
Agent-based approaches [8, 49, 16, 51, 58] hold great potential in long video understanding, because agents can simplify difficult problems and autonomously invoke a variety of tools to extract key information from long videos to assist in long video understanding. However, there is still significant room for improvement in current agent-based methods for long video tasks. Some methods use CLIP [39] to retrieve key frames for long video understanding [49, 58, 63, 55]. However, CLIP struggles to retrieve long-term temporal information and its pretrained data has a domain gap with long videos [46, 59]. Some approaches use external tools (e.g., memory bank [8, 13], RAG [21, 2, 16]) to aid MLLMs in long video understanding. Even with the support of tools, relying on a single MLLM to answer questions about long-form videos offers only a partial comprehension of the video content, leading to limited performance. In conclusion, there are two main challenges in long video understanding. The first challenge is how to better retrieve the video clip with key information according to the query. The second challenge is how to use multi-agent collaboration to better understand long videos.
@@
  • ␣Second, LVAgent is a novel multi-round and multi-step dynamical collaboration pipeline of MLLM Agents. LVAgent enables multi MLLM agents to process and integrate long-term temporal context dynamically for more comprehensive and robust long video understanding.
  • Finally, empirical results demonstrate the superior accuracy of our LVAgent. LVAgent achieves an accuracy exceeding 80% across four mainstream long video understanding tasks. It is the first multi-round dynamic multi-agent collaboration pipeline that outperforms all closed-source models (including GPT) and open-source models (including InternVL-2.5 and Qwen2-VL) in the long video understanding tasks. Notably, on the LongVideoBench dataset, LVAgent improves accuracy by up to 13.3% compared with SOTA.
@@ 3 Method
3.1 Selection
3.2 Perception
$D_{i}=A_{i}(\hat{v},Q,T,O)$ (2)
$K_{i}=A_{i}(\hat{v},Q,T,O)$ (3)
$K_{i}=A_{i}(H_{info},Q,T,O)$ (4)
$Sim(chk_{i})=ASP(f_{i},K_{i})+ASP(f_{i},Q)$ (5)
LongVR dataset for Long Video Retrieval.
To enhance the retrieval ability of ASP-CLIP to better assist with long video tasks, we compile a diverse dataset for fine-tuning the ASP-CLIP model, consisting of 82k video clips from five datasets, as shown in Table 8 in the appendix. This collection covers a wide range of scenes, themes, and video quality, providing a robust foundation for training a video retrieval model. We extract long videos from the datasets of ActivityNet [19], OpenVid-1M [36], ViTT [15], MovieChat-Caption [43], and Youcook2 [66] for the retrieval task, as depicted in Figure 6. First, we crop the videos clips with captions. If a video has been pre-annotated into multiple clips, we take these clips along with their captions as components of our dataset. Second, to better suit the long video scenario, we filter out video clips shorter than 5 seconds or longer than 12 minutes. Only clips of appropriate length and their captions are retained. Finally, considering that the context length of CLIP is limited, to prevent information truncation in the clip-text encoder due to overly long captions, we use Qwen2-VL [47] for caption optimization. Additionally, if the number of tokens in a caption is less than 20, we also filter out this data, as such short captions are insufficient to summarize the content of a relatively long video. After manually screening the captions, the resulting long video dataset has an average video duration of 145.6 seconds. The captions in the dataset have an average length of 71 tokens. In total, we assemble a long video retrieval dataset with 82K samples. The details of dataset constructions and data recipes can be found in the supplementary document. We finetune ASP-CLIP [3] on LongVR and achieved significant improvements.
@@ 3.2 Perception
Table 1: Comparison on Mainstream Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs". We are the first agent-based method to achieve over 80% accuracy on each long video understanding dataset
@@ 4.2 Implication Details
4.3 Comparison with SOTA
As shown in Table 1, we compare the methods with closed-source models [38, 44], open-source MLLMs [2, 24, 4, 47, 61, 40, 62, 22, 28, 20, 6], and agent-based systems [49, 51, 58, 33, 32]. Experimental results show that our method is the first agent-based method to achieve over 80% accuracy across these four benchmarks. Especially on LongVideoBench, LVAgent outperforms GPT-4o (SOTA) by 13.3%. On MLVU, LVAgent surpasses GPT-4o and InternVL-2.5 by 19.3% and 8.2%.
@@ 4.3 Comparison with SOTA
We evaluate LVAgent’s efficiency on the VideoMME benchmark, comparing it with state-of-the-art models in inference time and performance (Table 2). LVAgent processes 71.2 frames and 33.6s of inference time. Compared with the current MLLM models that perform best in long video tasks, our LVAgent can achieve results that significantly surpass these MLLMs while using fewer frames and less inference time.
@@ 4.4 Ablation Study
As shown in Table 4, we analyze the impact of agent combinations on LVBench and MLVU tasks using LLaVA-Video-72B (LV-72B), InternVL-2.5-8B (Intern-8B), and InternVL-2.5-78B (Intern-78B). Individually, LV-72B, Intern-8B, and Intern-78B achieve 68.3%, 67.8%, and 69.0% on LVBench, respectively. However, combining all three agents yields superior performance (80.0% on LVBench and 83.9% on MLVU). This improvement stems from the diversity of agents, enabling the model to leverage their complementary strengths for more robust reasoning.
@@
We investigate the effect of varying the number of maximum reflection rounds one, two, and three on multiple benchmarks, as illustrated in Figure 7. Increasing the number of rounds consistently enhances performance; for instance, LongVideoBench sees an increase from 69.0% to 80.0%. This improvement is attributed to the iterative refinement and consensus among agents, where multiple rounds progressively eliminate noisy knowledge and converge on more accurate answers, significantly boosting performance in complex reasoning.
@@
We further explore the impact of different retrieval models used in the Perception process, as shown in Table 5. We show the result of ASP-CLIP pretrained on our collected LongVR dataset. The results indicate that ASP-CLIP outperforms the other models.
@@ 6 Acknowledgement
We present a detailed breakdown of the data distribution and key statistics for LongVR dataset in Table 8. Specifically, ActivityNet-Caption contains 37,421 instances with an average caption length of 15.80 words and an average video duration of 171.44 seconds. OpenVid-1M comprises 30,000 instances, featuring longer captions with an average length of 143.16 words and shorter videos averaging 9.07 seconds. ViTT has 5,086 instances, with captions averaging 24.98 words and videos lasting 305.08 seconds on average. MovieChat-Caption includes 808 instances, having the longest captions at an average of 143.47 words, and videos with an average duration of 457.65 seconds. Youcook2 consists of 8,700 instances, with captions averaging 80.90 words and videos lasting 383.19 seconds. Collectively, the LongVR dataset encompasses a total of 82,015 instances, with an overall average caption length of 71.12 words and an average video duration of 145.62 seconds. This diverse and extensive dataset enables robust training of the ASP-CLIP model, capturing a wide range of video and caption characteristics.
@@
Appendix B More Experiment Details and Results.
B.1 Model Selection for Different Tasks
For VideoMME, we employ InternVL2.5 (78B & 8B) and Qwen2VL (72B). For EgoSchema, we use Qwen2VL (72B & 7B) and LLaVAVideo (72B). For MLVU and LongVideoBench, we utilize LLaVAVideo (72B) and InternVL2.5 (78B & 8B). Our selection mechanism chooses different agent groups based on pseudo labels and dataset-specific features.
B.2 Efficiency Analysis
We evaluate a total of 13,943 videos across the datasets LongVideoBench, MLVU, VideoMME (with and without subtitles), and EgoSchema. For comparison, Qwen2VL (72B) requires 350.5 hours in total, averaging 90.5 seconds per video. We break down the time consumption of our approach into the following steps: Fine-tuning, conducted with 95M parameters, takes 3 hours. Preselection on 750 videos consumes 6.58 hours, averaging 31.6 seconds per video. The evaluation phase takes 130.13 hours, with an average of 33.6 seconds per video. In total, the evaluation process (including preselection and evaluation) requires 139.71 hours, averaging 36.07 seconds per video—this is 2.5× faster than Qwen2VL.
B.3 The Result of Using More Than Three Agents
The performance of the 4-agent setup (EgoSchema: 83.0, LongVideoBench: 80.1, MLVU: 84.1, VideoMME: 81.8 / 86.5) is slightly superior to that of the 3-agent configuration, but it incurs additional time cost—taking 38.7 seconds per video, which is 5.1 seconds longer than the 3-agent setup. Considering the trade-off between accuracy and efficiency, we opt for the 3-agent configuration. Notably, LVAgent can be easily extended to incorporate more agents if needed.
Appendix C Prompts of LVAgent.